# DINO-WM: World Models on Pre-trained Visual Features enable Zero-shot Planning

### v1

@@ Abstract
The ability to predict future outcomes given control actions is fundamental for physical reasoning. However, such predictive models, often called world models, have proven challenging to learn and are typically developed for task-specific solutions with online policy learning. We argue that the true potential of world models lies in their ability to reason and plan across diverse problems using only passive data. Concretely, we require world models to have the following three properties: 1) be trainable on offline, pre-collected trajectories, 2) support test-time behavior optimization, and 3) facilitate task-agnostic reasoning. To realize this, we present DINO World Model (DINO-WM), a new method to model visual dynamics without reconstructing the visual world. DINO-WM leverages spatial patch features pre-trained with DINOv2, enabling it to learn from offline behavioral trajectories by predicting future patch features. This design allows DINO-WM to achieve observational goals through action sequence optimization, facilitating task-agnostic behavior planning by treating desired goal patch features as prediction targets. We evaluate DINO-WM across various domains, including maze navigation, tabletop pushing, and particle manipulation. Our experiments demonstrate that DINO-WM can generate zero-shot behavioral solutions at test time without relying on expert demonstrations, reward modeling, or pre-learned inverse models. Notably, DINO-WM exhibits strong generalization capabilities compared to prior state-of-the-art work, adapting to diverse task families such as arbitrarily configured mazes, push manipulation with varied object shapes, and multi-particle scenarios.
@@ 1 Introduction
Robotics and embodied AI has seen tremendous progress in recent years. Advances in imitation learning and reinforcement learning has enabled agents to learn complex behaviors across diverse tasks Lee et al. (2024); Zhao et al. (2023); Ma et al. (2024); Hafner et al. (2024); Hansen et al. (2024); Agarwal et al. (2022); Haldar et al. (2024). Despite this progress, generalization remains a major challenge Zhou et al. (2023). Existing approaches predominantly rely on policies that, once trained, operate in a feed-forward manner during deployment—mapping observations to actions without any further optimization or reasoning. Under this framework, successful generalization inherently requires agents to possess solutions to all possible tasks and scenarios once training is complete, which is only possible if the agent has seen similar scenarios during training Brohan et al. (2023b; a); Reed et al. (2022); Etukuru et al. (2024). However, it is neither feasible nor efficient to learn solutions for all potential tasks and environments in advance.
@@
Instead of learning the solutions to all possible tasks during training, an alternate is to fit a dynamics model on training data and optimize task-specific behavior during runtime. These dynamics models, also called world models Ha & Schmidhuber (2018), have a long history in robotics and control Sutton (1991); Todorov & Li (2005); Williams et al. (2017). More recently, several works have shown that world models can be trained on raw observational data Hafner et al. (2019; 2024); Micheli et al. (2023); Robine et al. (2023); Hansen et al. (2024). This enables flexible use of model-based optimization to obtain policies as it circumvents the need for explicit state-estimation. Despite this, significant challenges still remain in it use for solving general-purpose tasks.
@@
To understand the challenges in world modeling, let us consider the two broad paradigms in learning world models: online and offline. In the online setting, access to the environment is often required so data can be continuously collected to improve the world model, which in turn improves the policy and the subsequent data collection. However, the online world model is only accurate in the cover of the policy that was being optimized. Hence, while it can be used to train powerful task-specific policies , it requires retraining for every new task even in the same environment . Instead in the offline setting, the world model is trained on an offline dataset of collected trajectories in the environment, which removes its dependence on the task specificity given sufficient coverage in the dataset. However, when required to solve a task, methods in this domain require strong auxiliary information to overcome the lack of dense coverage on the task-specific domain. This auxiliary information can take the form of expert demonstrations Pathak et al. (2018), structured keypoints Ko et al. (2023); Wen et al. (2024), access to pretrained inverse models Du et al. (2023); Ko et al. (2023) or dense reward functions , all of which reduce the generality of using offline world models. The central question to building better offline world models is if there is alternate auxiliary information that does not compromise its generality?
@@
Figure 1: We present DINO-WM, a method for training visual models by using pretrained DINOv2 embeddings of image frames (a). Once trained, given a target observation $o_{T}$, we can directly optimize agent behavior by planning through DINO-WM using model-predictive control (b). The use of pretrained embeddings significantly improves performance over prior state-of-the-art world models (c).
@@
In this work, we present DINO-WM, a new and simple method to build task-agnostic world models from an offline dataset of trajectories. DINO-WM models the dynamics of world on compact embeddings of the world, rather than the raw observations themselves. For the embedding, we use pretrained patch-features from the DINOv2 model, which provides both a spatial and object-centric representation prior. We conjecture that this pretrained representation enables robust and consistent world modeling, which relaxes the necessity for task-specific data coverage. Given these visual embeddings and actions, DINO-WM uses the ViT architecture to predict future embeddings. Once this model is trained, planning to solve tasks is constructed as visual goal reaching, i.e. to reach a future desired goal given the current observation. Since the predictions by DINO-WM are high quality (see Figure 4), we can simply use model-predictive control with inference-time optimization to reach desired goals without any without any extra information during testing.
@@
DINO-WM is experimentally evaluated on four environment suites spanning maze navigation, sliding manipulation, and particle manipulation tasks. Our experiments reveal the following findings:
@@
Code and models for DINO-WM will be open-sourced to ensure reproducibility and videos of policies are made available on our project website: https://dino-wm.github.io.
@@ 2 Related Work
We build on top of several works in building world models, optimizing them, and using compact visual representations. For conciseness, we only discuss the ones most relevant to DINO-WM.
@@
Model-based Learning: Learning from models of dynamics has a rich literature spanning the fields of control, planning, and robotics Sutton (1991); Todorov & Li (2005); Astolfi et al. (2008); Holkar & Waghmare (2010); Williams et al. (2017). Recent works have shown that modeling dynamics and predicting future states can significantly enhance vision-based learning for embodied agents across various applications, including online reinforcement learning Hafner et al. (2024); Micheli et al. (2023); Hansen et al. (2024); Robine et al. (2023), exploration Mendonca et al. (2021; 2023a); Sekar et al. (2020), planning Finn & Levine (2017); Ebert et al. (2018); Hafner et al. (2019), and imitation learning Pathak et al. (2018). Several of these approaches initially focused on state-space dynamics Deisenroth & Rasmussen (2011); Chua et al. (2018); Lenz et al. (2015); Nagabandi et al. (2019), and has since been extended to handle image-based inputs, which we address in this work. These world models can predict future states in either pixel space Finn & Levine (2017); Ebert et al. (2018); Ko et al. (2023); Du et al. (2023) or latent representation space Yan et al. (2021). Predicting in pixel space, however, is computationally expensive due to the need for image reconstruction and the overhead of using diffusion models . On the other hand, latent-space prediction is typically tied to objectives of reconstructing images Hafner et al. (2019; 2024); Micheli et al. (2023), which raises concerns about whether the learned features contain sufficient information about the task. Moreover, many of these models incorporate reward prediction Hafner et al. (2024); Micheli et al. (2023); Robine et al. (2023), or use reward prediction as auxiliary objective to learn the latent representation Hansen et al. (2024; 2022), inherently making the world model task-specific. In this work, we aim to decouple task-dependent information from latent-space prediction, striving to develop a versatile and task-agnostic world model capable of generalizing across different scenarios.
@@
Generative Models as World Models: With the recent excitement of large scale foundation models, there have been initiatives on building large-scale video generation world models conditioned on agent’s actions in the domain of self-driving Hu et al. (2023), control Yang et al. (2023); Bruce et al. (2024), and general-purpose video generation Liu et al. (2024). These models aim to generate video predictions conditioned on text or high-level action sequences. While these models have demonstrated utility in downstream tasks like data augmentations, their reliance on language conditioning limits their application when precise visually indicative goals need to be reached. Additionally, the use of diffusion models for video generation makes them computationally expensive, further restricting their applicability for test-time optimization techniques such as MPC. In this work, we aim to build a world model in latent space rather than in the raw pixel space, which enables more precise planning and control.
@@
Pretrained Visual Representations: Significant advancements have been made in the field of visual representation learning, where compact features that capture spatial and semantic information can be readily used for downstream tasks. Pre-trained models like ImageNet pre-trained ResNet He et al. (2016), I-JEPA Assran et al. (2023), and DINO Caron et al. (2021); Oquab et al. (2024) for images, as well as V-JEPA Bardes et al. (2024) for videos, and R3M Nair et al. (2022), MVP Xiao et al. (2022) for robotics have allowed fast adaptation to downstream tasks as they contain rich spatial and semantic information. While many of these models represent images using a single global feature, the introduction of Vision Transformers (ViTs) Dosovitskiy et al. (2021) has enabled the use of pre-trained patch features, as demonstrated by DINO Caron et al. (2021); Oquab et al. (2024). DINO employs a self-distillation loss that allows the model to learn representations effectively, capturing semantic layouts and improving spatial understanding within images. In this work, we leverage DINOv2’s patch embeddings to train our world model, and demonstrate that it serves as a versatile encoder capable of handling multiple precise tasks.
@@ 3 DINO World Models
Overview and Problem formulation: Our work follows the vision-based control task framework, which models the environment as a partially observable Markov decision process (POMDP). The POMDP is defined by the tuple $(\mathcal{O},\mathcal{A},p)$, where $\mathcal{O}$ represents the observation space, and $\mathcal{A}$ denotes the action space. The environment’s dynamics are modeled by the transition distribution $p(o_{t+1}\mid o_{\leq t},a_{\leq t})$, which predicts future observations based on past actions and observations.
In this work, we aim to learn task-agonstic world models from pre-collected offline datasets, and use these world models to perform visual reasoning and control at test time. At test time, our system starts from an arbitrary environment state and is provided with a goal observation in the form of an RGB image, in line with prior works Wu et al. (2020); Ebert et al. (2018); Mendonca et al. (2023b), and is asked to perform a sequence of actions $a_{0},...,a_{T}$ such that the goal state can be achieved. This approach differs from world models used in online reinforcement learning (RL) where the objective is to optimize rewards for a fixed set of tasks at hand Hafner et al. (2024); Hansen et al. (2024), or from text-conditioned world models, where goals are specified through text prompts Du et al. (2023); Ko et al. (2023).
@@ 3.1 DINO-based World Models (DINO-WM)
Decoder model (optional for visualization): $\quad\hat{o}_{t}\sim q_{\theta}(o_{t}\mid z_{t})$
where the observation model encodes image observations to latent states $z_{t}$, and the transition model takes in a history of past latent states of length $H$. The decoder model takes in a latent $z_{t}$, and reconstruct the image observation $o_{t}$. We use $\theta$ to denote the parameters of these models. Note that our decoder is entirely optional, as the training objectives for the decoder is independent for training the rest part of the world model. This eliminates the need to reconstructing images both during training and testing, which reduces computational costs compared to otherwise coupling together the training of the observational model and the decoder, as in Hafner et al. (2024); Micheli et al. (2023).
DINO-WM models only the information available from offline trajectory data in an environment, in contrast to recent online RL world models that also require task-relevant information, such as rewards Hansen et al. (2022; 2024); Hafner et al. (2020), discount factors Hafner et al. (2022); Robine et al. (2023), and termination conditions Hafner et al. (2024); Micheli et al. (2023).
@@ 3.1.1 Observation Model
With the goal of learning a generic world model across many environments and the real world, we argue that the observation model should 1) be task and environment independent, and 2) contain rich spatial information which is crucial in navigation and manipulation tasks. Contrary to previous works where the observation model is always learned for the task at hand Hafner et al. (2024), we argue instead that it is not always possible for world models to learn an observation model from scratch when facing a new environment, as perception is a general task that can be learned from the large corpus of internet data. Therefore, we choose the out-of-the-box pre-trained DINOv2 model as our world model’s observation model, as it has been shown to excel at object detection, semantic segmentation, and depth estimation tasks which require substantial spatial understanding. The observation model is kept frozen throughout both training and testing time. At each time step $t$, it encodes an image $o_{t}$ to patch embeddings $z_{t}\in\mathbb{R}^{N\times E}$, where $N$ denotes the number of patches, and $E$ denotes the embedding dimension. This process is visualized in Figure 2.
@@ 3.1.2 Transition Model
We adopt the ViT Dosovitskiy et al. (2021) architecture for the transition model as it is a natural choice for processing patch features. However, a few modifications are required to the architecture to allow for additional conditioning on proprioception and controller actions.
Our transition model takes in a history of past latent states $z_{t-H:t-1}$ and actions $a_{t-H:t-1}$, where $H$ is a hyperparameter denoting the context length of the model, and predicts the latent state at next time step $z_{t}$. To properly capture the temporal dependencies, where the world state at time $t$ should only depend on previous observations and actions, we implement a causal attention mechanism in the ViT model, enabling the model to predict latents autoregressively at a frame level. Specifically, each patch vector $z^{i}_{t}$ for the latent state $z_{t}$ attends to $\{z^{i}_{t-H:t-1}\}_{i=1}^{N}$. This is different from past work IRIS Micheli et al. (2023) which similarly represent each observation as a sequence of vectors, but autoregressively predict $z^{i}_{t}$ at a token level, attending to $\{z^{i}_{t-H:t-1}\}_{i=1}^{N}$ as well as $\{z^{i}_{t}\}_{i=1}^{<k}$. We argue that predicting at a frame level and treating patch vectors of one observation as a whole better captures global structure and temporal dynamics, modeling dependencies across the entire observation rather than isolated tokens, leading to improved temporal generalization.

### v2

@@ Abstract
The ability to predict future outcomes given control actions is fundamental for physical reasoning. However, such predictive models, often called world models, remains challenging to learn and are typically developed for task-specific solutions with online policy learning. To unlock world models’ true potential, we argue that they should 1) be trainable on offline, pre-collected trajectories, 2) support test-time behavior optimization, and 3) facilitate task-agnostic reasoning. To this end, we present DINO World Model (DINO-WM), a new method to model visual dynamics without reconstructing the visual world. DINO-WM leverages spatial patch features pre-trained with DINOv2, enabling it to learn from offline behavioral trajectories by predicting future patch features. This allows DINO-WM to achieve observational goals through action sequence optimization, facilitating task-agnostic planning by treating goal features as prediction targets. We demonstrate that DINO-WM achieves zero-shot behavioral solutions at test time on six environments without expert demonstrations, reward modeling, or pre-learned inverse models, outperforming prior state-of-the-art work across diverse task families such as arbitrarily configured mazes, push manipulation with varied object shapes, and multi-particle scenarios.
Machine Learning, ICML
@@ 1 Introduction
Robotics and embodied AI have seen tremendous progress in recent years. Advances in imitation learning and reinforcement learning have enabled agents to learn complex behaviors across diverse tasks (Agarwal et al., 2022; Zhao et al., 2023; Lee et al., 2024; Ma et al., 2024; Hafner et al., 2024; Hansen et al., 2024; Haldar et al., 2024; Jia et al., 2024). Despite this progress, generalization remains a major challenge (Zhou et al., 2023). Existing approaches predominantly rely on policies that, once trained, operate in a feed-forward manner during deployment—mapping observations to actions without any further optimization or reasoning. Under this framework, successful generalization inherently requires agents to possess solutions to all possible tasks and scenarios once training is complete, which is only possible if the agent has seen similar scenarios during training (Reed et al., 2022; Brohan et al., 2023b, a; Etukuru et al., 2024). However, it is neither feasible nor efficient to learn solutions for all potential tasks and environments in advance.
@@
Instead of learning the solutions to all possible tasks during training, an alternate is to fit a dynamics model on training data and optimize task-specific behavior at runtime. These dynamics models, also called world models (Ha & Schmidhuber, 2018), have a long history in robotics and control (Sutton, 1991; Todorov & Li, 2005; Williams et al., 2017). More recently, several works have shown that world models can be trained on raw sensory data (Hafner et al., 2019; Micheli et al., 2023; Robine et al., 2023; Hansen et al., 2024; Hafner et al., 2024). This enables flexible use of model-based optimization to obtain policies as it circumvents the need for explicit state-estimation. Despite this, significant challenges remains in its use for solving general-purpose tasks.
@@
To understand the challenges in world modeling, let us consider the two broad paradigms in learning world models: online and offline. In the online setting, access to the environment is often required so data can be continuously collected to improve the world model, which in turn improves the policy and the subsequent data collection. However, the online world model is only accurate in the cover of the policy that was being optimized. Hence, while it can be used to train powerful task-specific policies, it requires retraining for every new task even in the same environment. Instead, in the offline setting, the world model is trained on an offline dataset of collected trajectories in the environment, which removes its dependence on the task specificity given sufficient coverage in the dataset. However, when required to solve a task, methods in this domain require strong auxiliary information which can take the form of expert demonstrations (Pathak et al., 2018; Wang et al., 2023), structured keypoints (Ko et al., 2023; Wen et al., 2024), access to pretrained inverse models (Du et al., 2023; Ko et al., 2023) or dense reward functions (Ding et al., 2024), all of which reduce the generality of using offline world models. The central question to building better offline world models is if there is alternate auxiliary information that does not compromise its generality?
@@
Figure 1: We present DINO-WM, a method for training visual models by using pretrained DINOv2 embeddings of image frames (a). Once trained, given a target observation $o_{T}$, we can directly optimize agent behavior by planning through DINO-WM using model predictive control (b). The use of pretrained embeddings significantly improves performance over prior state-of-the-art world models (c).
@@
In this work, we present DINO-WM, a new and simple method to build task-agnostic world models from an offline dataset of trajectories. DINO-WM models the world dynamics on compact embeddings of the world, rather than the raw observations themselves. For the embedding, we use pretrained patch-features from the DINOv2 model, which provides both a spatial and object-centric representation prior. We conjecture that this pretrained representation enables robust and consistent world modeling, which relaxes the necessity for task-specific data coverage. Given these visual embeddings and actions, DINO-WM uses the ViT architecture to predict future embeddings. Once this model is trained on the offline dataset, planning to solve tasks is constructed as visual goal reaching, i.e. to reach a future desired goal given the current observation. Since the predictions by DINO-WM are high quality (see Figure 4), we can simply use model predictive control with inference-time optimization to reach desired goals without any extra information during testing.
@@
DINO-WM is experimentally evaluated on six environment suites spanning maze navigation, sliding manipulation, robotic arm control, and deformable object manipulation tasks. Our experiments reveal the following findings:
@@
Code and models for DINO-WM are open-sourced to ensure reproducibility and videos of planning are made available on our anonymous project website: https://dino-wm.github.io.
@@ 2 Related Work
We build on top of several works in developing world models, optimizing behaviors from them, and leveraging compact visual representations. For conciseness, we only discuss the ones most relevant to DINO-WM.
@@
Model-based Learning: Learning from models of dynamics has a rich literature spanning the fields of control, planning, and robotics (Sutton, 1991; Todorov & Li, 2005; Astolfi et al., 2008; Holkar & Waghmare, 2010; Williams et al., 2017). Recent works have shown that modeling dynamics and predicting future states can significantly enhance vision-based learning for embodied agents across various applications, including online reinforcement learning (Micheli et al., 2023; Robine et al., 2023; Hansen et al., 2024; Hafner et al., 2024), exploration (Sekar et al., 2020; Mendonca et al., 2021, 2023a), planning (Watter et al., 2015; Finn & Levine, 2017; Ebert et al., 2018; Hafner et al., 2019), and imitation learning (Pathak et al., 2018). Several of these approaches initially focused on state-space dynamics (Deisenroth & Rasmussen, 2011; Lenz et al., 2015; Chua et al., 2018; Nagabandi et al., 2019), and have since been extended to handle image-based inputs, which we address in this work. These world models can predict future states in either pixel space (Finn & Levine, 2017; Ebert et al., 2018; Ko et al., 2023; Du et al., 2023) or latent representation space (Yan et al., 2021). Predicting in pixel space, however, is computationally expensive due to the need for image reconstruction and the overhead of using diffusion models (Ko et al., 2023). On the other hand, latent-space prediction is typically tied to objectives of reconstructing images (Hafner et al., 2019; Micheli et al., 2023; Hafner et al., 2024), which raises concerns about whether the learned features contain sufficient information about the task. Moreover, many of these models incorporate reward prediction (Micheli et al., 2023; Robine et al., 2023; Hafner et al., 2024), or use reward prediction as auxiliary objective to learn the latent representation (Hansen et al., 2022, 2024), inherently making the world model task-specific. In this work, we aim to decouple task-dependent information from latent-space prediction, striving to develop a versatile and task-agnostic world model capable of generalizing across different scenarios.
@@
Generative Models as World Models: With the recent excitement of large scale foundation models, there have been initiatives on building large-scale video generation world models conditioned on agent’s actions in the domain of self-driving (Hu et al., 2023), control (Yang et al., 2023; Bruce et al., 2024), and general-purpose video generation (Liu et al., 2024). These models aim to generate video predictions conditioned on text or high-level action sequences. While these models have demonstrated utility in downstream tasks like data augmentations, their reliance on language conditioning limits their application when precise visually indicative goals need to be reached. Additionally, the use of diffusion models for video generation makes them computationally expensive, further restricting their applicability for test-time optimization techniques such as MPC. In this work, we aim to build a world model in latent space instead of raw pixel space, enabling more precise planning and control.
@@
Pretrained Visual Representations: Significant advancements have been made in the field of visual representation learning, where compact features that capture spatial and semantic information can be readily used for downstream tasks. Pre-trained models like ImageNet pre-trained ResNet (He et al., 2016), I-JEPA (Assran et al., 2023), and DINO (Caron et al., 2021; Oquab et al., 2024) for images, as well as V-JEPA (Bardes et al., 2024) for videos, and R3M (Nair et al., 2022), MVP (Xiao et al., 2022) for robotics have allowed fast adaptation to downstream tasks as they contain rich spatial and semantic information. While many of these models represent images using a single global feature, the introduction of Vision Transformers (ViTs) (Dosovitskiy et al., 2021) has enabled the use of pre-trained patch features, as demonstrated by DINO (Caron et al., 2021; Oquab et al., 2024). DINO employs a self-distillation loss that allows the model to learn representations effectively, capturing semantic layouts and improving spatial understanding within images. In this work, we leverage DINOv2’s patch embeddings to train our world model, and demonstrate that it serves as a versatile encoder capable of handling various precise tasks.
@@ 3 DINO World Models
Overview and Problem formulation: Our work follows the vision-based control task framework, which models the environment as a partially observable Markov decision process (POMDP). The POMDP is defined by the tuple $(\mathcal{O},\mathcal{A},p)$, where $\mathcal{O}$ represents the observation space, and $\mathcal{A}$ denotes the action space. The dynamics of the environment is modeled by the transition distribution $p(o_{t+1}\mid o_{\leq t},a_{\leq t})$, which predicts future observations based on past actions and observations.
In this work, we aim to learn task-agnostic world models from precollected offline datasets, and use these world models to perform visual reasoning and control at test time. At test time, our system starts from an arbitrary environment state and is provided with a goal observation in the form of an RGB image, in line with prior works (Ebert et al., 2018; Wu et al., 2020; Mendonca et al., 2023b), and is asked to perform a sequence of actions $a_{0},...,a_{T}$ to reach the goal state. This approach differs from the world models used in online reinforcement learning (RL) where the objective is to optimize the rewards for a fixed set of tasks at hand (Hafner et al., 2024; Hansen et al., 2024), or from text-conditioned world models, where the goals are specified through text prompts (Du et al., 2023; Ko et al., 2023).
@@ 3.1 DINO-based World Models (DINO-WM)
Decoder model: $\quad\hat{o}_{t}\sim q_{\theta}(o_{t}\mid z_{t})$
(optional for visualization)
where the observation model encodes image observations to latent states $z_{t}$, and the transition model takes in a history of past latent states of length $H$. The decoder model takes in a latent $z_{t}$, and reconstructs the image observation $o_{t}$. We use $\theta$ to denote the parameters of these models. Note that our decoder is entirely optional, as the training objectives for the decoder is independent for training the rest part of the world model. This eliminates the need to reconstruct images both during training and testing, which reduces computational costs compared to otherwise coupling together the training of the observational model and the decoder, as in (Micheli et al., 2023; Hafner et al., 2024). We ablate and show the effectiveness of this choice in Appendix A.4.3.
DINO-WM models only the information available from offline trajectory data in an environment, in contrast to recent online RL world models that also require task-relevant information, such as rewards (Hafner et al., 2020; Hansen et al., 2022, 2024), discount factors (Hafner et al., 2022; Robine et al., 2023), and termination conditions (Micheli et al., 2023; Hafner et al., 2024).
@@ 3.1.1 Observation Model
To learn a generic world model across many environments and the real world, we argue that the observation model should 1) be task and environment independent, and 2) capture rich spatial information for navigation and manipulation. Contrary to previous works where the observation model is always learned for the task at hand (Hafner et al., 2024), we argue instead that it can be inefficient and often not possible to learn a good observation model from scratch when facing a new environment, as perception is a general task that benefits from large-scale internet data. Therefore, we use the pre-trained DINOv2 model as our world model’s observation model, leveraging its strong spatial understanding for tasks like object detection, semantic segmentation, and depth estimation (Oquab et al., 2024). The observation model remains frozen during training and testing. At each time step $t$, it encodes an image $o_{t}$ to patch embeddings $z_{t}\in\mathbb{R}^{N\times E}$, where $N$ denotes the number of patches, and $E$ denotes the embedding dimension. This process is visualized in Figure 2.
@@ 3.1.2 Transition Model
We adopt the ViT architecture (Dosovitskiy et al., 2021) for the transition model due to its suitability for processing patch features. We remove the tokenization layer, as it operates on patch embeddings, effectively transforming it into a decoder-only transformer. We further make a few modifications to the architecture to allow for additional conditioning on proprioception and controller actions.
Our transition model takes in a history of past latent states $z_{t-H:t-1}$ and actions $a_{t-H:t-1}$, where $H$ is a hyperparameter denoting the context length of the model, and predicts the latent state at next time step $z_{t}$. To properly capture the temporal dependencies, where the world state at time $t$ should only depend on previous observations and actions, we implement a causal attention mechanism in the ViT model, enabling the model to predict latents autoregressively at a frame level. Specifically, each patch vector $z^{i}_{t}$ for the latent state $z_{t}$ attends to $\{z^{i}_{t-H:t-1}\}_{i=1}^{N}$. This is different from past work IRIS (Micheli et al., 2023) which similarly represents each observation as a sequence of vectors, but autoregressively predict $z^{i}_{t}$ at a token level, attending to $\{z^{i}_{t-H:t-1}\}_{i=1}^{N}$ as well as $\{z^{i}_{t}\}_{i=1}^{<k}$. We argue that predicting at a frame level and treating patch vectors of one observation as a whole better captures global structure and temporal dynamics, modeling dependencies across the entire observation rather than isolated tokens, leading to improved temporal generalization. The effectiveness of this attention mask has been shown in our ablation experiments in Appendix A.4.2